# Mechanical Optimization of Skateboard Pumping

### 29 Dec 2023 v1

@@ Abstract
Skateboarders perform a reciprocating motion on a curved ramp, called pumping, by moving their bodies up and down perpendicular to the ramp’s surface. We propose a simple mechanical model for this pumping motion and solve the equation of motion explicitly in angular coordinates. This allows us to derive an optimal control strategy to maximize amplitude by dynamically adjusting the center of mass of the skateboarder. This optimal strategy is compared to experimental results for the motion of a skilled and an unskilled skateboarder in a half-pipe, validating that a skilled skateboarder follows the optimal control strategy more closely.
@@
where $\theta$ is the angle relative to $y$-axis, $L$ is the radius of the cylindrical ramp, $h$ is the height of the center of mass, and $g\sin\theta$ is the gravitational restoring force. A skateboarder controls his/her center of mass with respect to height, $h=h(t)$, perpendicular to the surface of the ramp. For more details on the dynamics of swinging, we refer to recent works on the pumping of a swing taking into account the detailed aspects of swinging specific to human body movement [19, 20, 21, 22]. Due to the conservation of angular momentum in the frictionless regime, a change in body height at just the right moment will result in a gain in rotational motion. We therefore regard the height as an external control function to be optimised for maximal amplitude gain under certain constraints. Firstly, we assume that the height is constrained between a minimal bending and a maximal standing posture, $H_{0}\leq h\leq H_{1}$. Secondly, we assume that the energy input through the muscular activity of the human or the control force of a skateboarding robot is limited, leading to a constraint in the acceleration of the height $|\ddot{h}|\leq g$.
We consider equation (1) together with kinetic friction by assuming that the wheels of the skateboard kinetic viscosity $\mu N$ at the surface of the ramp, where $\mu$ is a material-specific friction coefficient and $N$ is the normal force
@@
Starting from an initial angle $\theta_{0}$ and zero initial velocity $\dot{\theta}_{0}=0$, the height $h$ is adjusted during a role from the initial angle to the final angle $\theta^{*}$ at which all kinetic energy is consumed and the amplitude is maximal. Different height variations then lead to different maximal amplitudes $\theta^{*}$. We are interested in an optimal control function $\hat{h}=\hat{h}(t)$ such that
@@
Figure 2: Pumping on a cylindrical ramp in the presence of kinetic friction. (a) The attained angle $\theta^{*}$ plotted against the pumping timing $\delta$ and pumping momentum $\beta$. (b) The optimal control height $h$ and its acceleration $\ddot{h}$ in the frictionless case (solid black line); unbounded control in the presence of kinetic friction (green dashed line) and the optimal solution under the acceleration bounded by $|\ddot{h}|\leq g$ (thick blue line). The optimal strategies for a cylindrical surface are compared with unsuccessful controls where an achieved angle $\theta^{*}$ is smaller than the initial angle $|\theta_{0}|$ (brown and orange dotted lines). Parameters: $\theta_{0}=-\pi/4\approx-0.79,L=3,H_{0}=0.9,H_{1}=1.0$ and $\mu=0.02$.
To tackle the optimization problem (3) directly, we assume that the control height is a monotonic function of time for a one-directional roll along the cylindrical ramp. Since the angle $\theta(t)$ along one roll is a monotonic function of time as well, we may rewrite the governing equation for the radius $r(\theta)=L-h(t)$ as a function of the angle, which will simplify the notation in the following. This allows us to derive an equation for the kinetic energy with respect to angular rotation $E_{\rm kin}=\dot{\theta}^{2}/2$ as a function of $\theta$. The analytical trick to write the kinetic energy and the radius as functions of the angle - the original dynamical variable - is reminiscent of the modern solution to the Brachistochrome problem [25] and similar problems in the calculus of variations.
@@
We remark that the influence of an air drag or an inertial drag proportional to the square of the velocity, $F_{\rm air}=C_{D}h|\mathbf{v}|\mathbf{v}$, for some dimensionless drag coefficient $C_{D}$, can be treated analogously to kinetic friction and an explicit formula similar to (5) may be derived by explicitly solving a linear differential equation for the angular kinetic energy. The numerical results for the optimal control strategy, however, show little deviation as compared to pure kinetic friction and we focus on formula (5) henceforth. At the maximal amplitude $\theta^{*}$ along one monotonic roll, all kinetic energy is consumed and we have that $E_{\rm kin}(\theta^{*})=0$. Setting $\mu=0$ in (5) we immediately see that the optimal control strategy in the frictionless case is given by
@@
$r(\theta)=L-H_{0}-\frac{H_{1}-H_{0}}{1+e^{\beta(\theta-\delta)}},$ (8)
@@
For the experimental validation of the optimal control strategy, two skateboarders with different degrees of skill were asked to get to the top of the ramp as fast as possible by adopting their most efficient pumping action. The skilled skateboarder has eleven years of experience, while the unskilled skateboarder has two years of experience. For each trial, they started from a position of rest and stopped their pumping when they reached the top of the ramp. The subject’s center of mass was estimated automatically by the motion-capture software Xsens. The acceleration in the sagittal plane $\{a_{x},a_{y}\}$ was translated into the height acceleration measured relative to the ramp surface $\ddot{h}$ using information about the center of mass of the subjects.
@@
We conclude with a summary of the results. We modeled the pumping motion of a skateboarder in a half-pipe as a variable-length pendulum with kinetic friction in the cylindrical part of the ramp’s surface and solved the equation of motion explicitly for the kinetic energy in angular coordinates. This allows us to formulate an optimization problem for the maximal angle in one roll. The results were compared to an experimental study in which the pumping strategies of a skilled and an unskilled skateboarder trying to gain height as fast as possible were compared. The skilled skateboarder showed less deviation from the theoretically predicted optimal control strategy.
By simplifying such the highly complex and delicate movement of skateboarding into the low-degree-of-freedom dynamical problem, we were able to suggest potential improvements in the performance of an athlete. This type of simplified kinematic consideration could also be applied to investigate improvements in other sports, such as ski jumping.
@@ Acknowledgements.
We thank Kazuo Tsuchiya and Eiji Uchibe for constructive comments, and Kai Shinomoto for drawing the illustrations for Figure 3. D.E.C. and S.S. were supported by the New Energy and Industrial Technology Development Organization (NEDO). F.K. was supported by the European Research Council (ERC) Advanced Grant 834763-PonD.

### 16 Mar 2024 v2

@@ Abstract
Skateboarders perform a reciprocating motion on a curved ramp, called pumping, by moving their bodies up and down perpendicular to the ramp surface. We propose a simple mechanical model for this pumping motion and solve the equation of motion explicitly in angular coordinates. This allows us to derive an optimal control strategy to maximize amplitude by dynamically adjusting the center of mass of the skateboarder. This optimal strategy is compared to experimental results for the motion of a skilled and an unskilled skateboarder in a half-pipe, validating that a skilled skateboarder follows the optimal control strategy more closely.
@@
where $\theta$ is the angle relative to a vertical axis, $L$ is the radius of the cylindrical ramp, $h$ is the height of the center of mass, and $g\sin\theta$ is the gravitational restoring force. A skateboarder controls his/her center of mass with respect to height, $h=h(t)$, perpendicular to the surface of the ramp. For more details on the dynamics of swinging, we refer to recent works on the pumping of a swing taking into account the detailed aspects of swinging specific to human body movement [19, 20, 21, 22]. The mechanism in which a change in body height will result in a gain in rotational motion may be intuitively understood as a result of the conservation of angular momentum, particularly when the friction can be ignored. We therefore regard the height as an external control function to be optimised for maximal amplitude gain under certain constraints. Firstly, we assume that the height is constrained between a minimal bending and a maximal standing posture, $H_{0}\leq h\leq H_{1}$. Secondly, we assume that the energy input through the muscular activity of the human or the control force of a skateboarding robot is limited, leading to a constraint in the acceleration of the height $|\ddot{h}|\leq g$.
We consider equation (1) together with kinetic friction by assuming that the wheels of the skateboard kinetic friction $\mu N$ at the surface of the ramp, where $\mu$ is a material-specific friction coefficient and $N$ is the normal force
@@
Starting from an initial angle $\theta_{0}$ and zero initial velocity $\dot{\theta}_{0}=0$, the height $h$ is adjusted during a roll from the initial angle to the final angle $\theta^{*}$ at which all kinetic energy is consumed and the amplitude is maximal. Different height variations then lead to different maximal amplitudes $\theta^{*}$. We are interested in an optimal control function $\hat{h}=\hat{h}(t)$ such that
@@
Figure 2: Pumping on a cylindrical ramp in the presence of kinetic friction. (a) The attained angle $\theta^{*}$ plotted against the pumping timing $\delta$ and pumping momentum $\beta$. (b) The optimal height $h$ and its acceleration $\ddot{h}$ in the frictionless case (solid black line); unbounded control in the presence of kinetic friction (green dashed line) and the optimal solution under the acceleration bounded by $|\ddot{h}|\leq g$ (thick blue line). The optimal strategies for a cylindrical ramp are compared with unsuccessful controls where an achieved angle $\theta^{*}$ is smaller than the initial angle $|\theta_{0}|$ (brown and orange dotted lines). Parameters: $\theta_{0}=-\pi/4\approx-0.79,L=3,H_{0}=0.9,H_{1}=1.0$ and $\mu=0.02$.
To tackle the optimization problem (3) directly, we assume that the height is a monotonic function of time for a one-directional roll along the cylindrical ramp. Since the angle $\theta(t)$ along one roll is a monotonic function of time as well, we may rewrite the governing equation for the radius $r(\theta)=L-h(t)$ as a function of the angle, which will simplify the notation in the following. This allows us to derive an equation for the kinetic energy with respect to angular rotation $E_{\rm kin}=\dot{\theta}^{2}/2$ as a function of $\theta$. The analytical trick to write the kinetic energy and the radius as functions of the angle - the original dynamical variable - is reminiscent of the modern solution to the Brachistochrome problem [25] and similar problems in the calculus of variations.
@@
We remark that the influence of an air drag or an inertial drag proportional to the square of the velocity, $F_{\rm air}=C_{D}|\mathbf{v}|\mathbf{v}$ can be treated analogously to kinetic friction and an explicit formula similar to (5) may be derived by explicitly solving a linear differential equation for the angular kinetic energy. The numerical results for the optimal control strategy, however, show little deviation as compared to pure kinetic friction and we focus on formula (5) henceforth. At the maximal amplitude $\theta^{*}$ along one monotonic roll, all kinetic energy is consumed and we have that $E_{\rm kin}(\theta^{*})=0$. Setting $\mu=0$ in (5) we immediately see that the optimal control strategy in the frictionless case is given by
@@
$r(\theta)=L-H_{0}-\frac{H_{1}-H_{0}}{1+e^{-\beta(\theta-\delta)}},$ (8)
@@
For the experimental validation of the optimal control strategy, two skateboarders with different degrees of skill were asked to gain height as much as possible by adopting their most efficient pumping action. The skilled skateboarder has eleven years of experience, while the unskilled skateboarder has two years of experience. For each trial, they started from a position of rest and stopped their pumping when they reached the top of the ramp. The subject’s center of mass was estimated automatically by the motion-capture software Xsens. The acceleration in the sagittal plane $\{a_{x},a_{y}\}$ was translated into the height acceleration measured relative to the ramp surface $\ddot{h}$ using information about the center of mass of the subjects.
@@
We conclude with a summary of the results. We modeled the pumping motion of a skateboarder in a half-pipe as a variable-length pendulum with kinetic friction in the cylindrical part of the ramp surface and solved the equation of motion explicitly for the kinetic energy in angular coordinates. This allows us to formulate an optimization problem for the maximal angle in one roll. The results were compared to an experimental study in which the pumping strategies of skilled and unskilled skateboarders trying to gain height were compared. The skilled skateboarder showed less deviation from the theoretically predicted optimal control strategy.
By simplifying the highly complex and delicate movement of skateboarding into the low-degree-of-freedom dynamical problem, we were able to suggest potential improvements in the performance of an athlete. This type of simplified kinematic consideration could also be applied to investigate improvements in other sports, such as ski jumping.
@@ Acknowledgements.
We thank Kazuo Tsuchiya, Eiji Uchibe, and Shin Takagi for constructive comments, and Kai Shinomoto for drawing the illustrations for Figure 3. D.E.C. and S.S. were supported by the New Energy and Industrial Technology Development Organization (NEDO).